help me solve this set of linear equations
5x-y=3
y=-3x-11
thank you very much

Answer by nerdybill(7384)

5x-y=3
y=-3x-11
.
substitute the value of y from 2nd equation into equation 1 and solve for x:
starting with:
5x-y=3
apply substitution:
5x-(-3x-11)=3
5x+3x+11=3
8x+11=3
8x = -8
x = -1
.
substitute above into equation 2 and solve for y:
y=-3x-11
y=-3(-1)-11
y=3-11
y = -8
.
answer: (x,y) = (-1, -8)